A ceramic jar contains 48 quarters and 21 dimes. If two coins are randomly drawn in succession without replacement what is the probability that the total value of the coins is .20? I have taken 21/69 x 20/68 amd come up with the answer of 5/391 but this is not one of the options. The answers given to me on the study sheet are 188/391 147/1564 35/391 or 84/391. I am not sure what I have done wrong if anything and how this number is achieved. Help please
Answer by stanbon(75887)

A ceramic jar contains 48 quarters and 21 dimes. If two coins are randomly drawn in succession without replacement what is the probability that the total value of the coins is .20? I have taken 21/69 x 20/68 amd come up with the answer of 5/391 but this is not one of the options. The answers given to me on the study sheet are 188/391 147/1564 35/391 or 84/391. I am not sure what I have done wrong if anything and how this number is achieved. Help please
------
Your procedure in correct.
Your fractions are correct.
Somehow a mistanke was made when you multiplied (21/69)(20/68)
# of ways to succeed: 21C2 = 210
# of random pairs: 69C2 = 2346
---
P(2 dimes) = 210/2346 = 35/391
==================================
Cheers,
Stan H.
=====================